Reports 1-1 of 1 Clear search Modify search
MIF (General)
takafumi.ushiba - 20:17 Friday 04 March 2022 (19983) Print this report
Contnuous lock test of GRX

I kept locking X-arm with GRX by guardian in this weekend.
So, if someone want to use PR3, ITMX, and ETMX, please request DOWN state to XARM guardian, first.

In addition, after finishing the work, please move back the suspensions to ALIGNED state (or comissioning mode for PR3, which is written in klog19674).
Then, request ALS_LOCKED state to XARM guardian and recover the Xarm lock.

Comments to this report:
takafumi.ushiba - 11:44 Monday 07 March 2022 (19990) Print this report

I stopped the continuous X-arm lock at 8:55 JST.

I found many short term lock loss (~10 ms), which cannot be detected by the EPICS channels like GRD-XARM_STATE_N (fig1).
10 ms is too short to follow the suspension motion, so it is likely that there are some causes of the lock loss except for the suspension motion.

Images attached to this comment
takafumi.ushiba - 9:00 Tuesday 08 March 2022 (20005) Print this report

I continued to lock X arm with GR from 2022-03-07 21:00 JST to 2022-03-08 9:00 JST.

takafumi.ushiba - 19:00 Tuesday 08 March 2022 (20013) Print this report

[Yamamoto, Ushiba]

Yamamoto-kun found the one possible reason of the unstability of the lock, which is reported in klog6638.
Current offset to VCO was set as 0.5 V while it should be between 0.8 V to 2.8 V.
So, we changed the offfset and set it as 1.8 V.

After the change, it seems much more stable but still have a glitch like lock loss.
We need to continue the investigation.

takahiro.yamamoto - 22:11 Tuesday 08 March 2022 (20015) Print this report
Output voltage to VCO is related to the stability of green lock. But a offset change doesn't seem to become a final solution for this problem.

Fig.1 shows the latest situation of green lock and green curve represents a output to the VCO. Due to the long-term drift (~a few tens of minutes) output voltage goes out of linear range.
Same behavior can be seen other time. Fig.2 shows the PDHX signals in this morning. It's before changing offset value. But lock is stable in a few hours because output voltage is in the linear range for a long time.
It seems to correlate with the drift of arm input signal. (We need to check more carefully.)

Seeing signal with short time as shown in Fig.3, 1Hz (it should comes from Type-C resonance) is dominant in RMS. By locking Green-X without IMC lock, local problem in ALS may be able to be isolated from ones in others.
Images attached to this comment
Search Help